simultaneous equations
2x-4y=8
x-2y=4

Answer by checkley77(12844)

2x-4y=8
x-2y=4 multiply this equation by -2 & add
-2x+4y=-8
2x-4y=8
--------------------
0x+0y=0 no unique solution for x or y.
These are 2 identical equations or parallel lines if graphed.